If m = 2, find the value of:
3m – 5....................

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that m = 2
Then, substitute the value of p in the question
= (3 × 2)–5
= 6–5
= 1

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Gven : m = 2
Substituting value of m in the equation,
3m - 5 = 3×2 - 5 = 6-5 = 1
 
Hope it helps!
Thank you ✌️